OpenStudy (anonymous):

s − 4 = square root of 8s + 16

OpenStudy (anonymous):

(s-4)^2 = 8s + 16 s^2 -8s +16 = 8s + 16 s^2 = 16s s = 16

OpenStudy (anonymous):

thanks